𝑓(π‘₯) = 𝑒 βˆ’π‘₯ 1+π‘’βˆ’π‘₯ . i) Determine whether 𝑓(π‘₯) is a one-to-one function.Β 


1
Expert's answer
2021-07-27T10:57:10-0400

Let's solve the equation "f(x)=f(y)" :

"\frac{e^{-x}}{1+e^{-x}}=\frac{e^{-y}}{1+e^{-y}}"

"\frac{1}{e^x+1}=\frac{1}{e^y+1}"

"e^x+1=e^y+1"

"e^x=e^y"

"x=y"

This means that the function "f(x)" is one-to-one.